OpenStudy (anonymous):

multiply (6 +sqrt2)(6-sqrt 2)

OpenStudy (anonymous):

36-6sqrt2+6sqrt2-2=36-2=34

OpenStudy (anonymous):

6*6 =36 +sqrt2 * -sqrt2 = -2 6 * -sqrt2 cancels out 6 * + sqrt 2 so 36-2 = 34

OpenStudy (anonymous):

=36-2=34 ( using idendity of 2 square difference)